OpenStudy (anonymous):

Write a quadratic equation in the variable x having the given numbers as solutions. Type the equation in standard form , ax^2+bx+c=0. The equation is _=0. Please help.

OpenStudy (anonymous):

Please, could you write the equation again?

OpenStudy (anonymous):

ax^2+bx+c=0 Solution is 4 only The equation is_=0

OpenStudy (anonymous):

I think I don't understand your question. Maybe this is useful: \[x=(-b \pm \sqrt{b ^{2}-4ac})/(2a)\]

OpenStudy (anonymous):

oh, got it. It's this way: (x-4)^2=x^2-8x+16=0

OpenStudy (anonymous):

Is that useful?

OpenStudy (anonymous):

Yes, that was helpful for I had no clue about this one. Thank you

OpenStudy (anonymous):

You are welcome!